OpenStudy (anonymous):

Find the square of 25/9 A. 5/3 B. 3/5 C. 7 58/81 D. 15/2 E. 650/81

OpenStudy (abb0t):

\[\frac{ 5 \times 5 }{ 3 \times 3 } = \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

The square or the square root?

OpenStudy (anonymous):

square

OpenStudy (anonymous):

The square is going to be \[ \frac{25}{9}\times \frac{25}{9} \]

OpenStudy (jiteshmeghwal9):

i think his question is wrong since his options are right when the question is asked for square root .

OpenStudy (jiteshmeghwal9):

so,